A sample of Helium has a volume of 2 Liters when the pressure is 250 torr, What volume does the gas occupy at 100 torr?
1
Expert's answer
2020-10-22T08:38:54-0400

At constant temperature, the volume "V" of an ideal gas is inversely proportional to its pressure "p":

"V\propto\frac{1}{p}" , or "p_1V_1 = p_2V_2"

where "p_1" and "V_1" are pressure and volume in the initial state , respectively, and "p_2" and "V_2" are pressure and volume in the final state , respectively.


In the initial state, the system has pressure "p_1" of 250 torr and its volume "V_1" is 2 L. In the final state, the pressure "p_2" is 100 torr. Using the relation above, we get the volume in the final state:

"V_2 = \frac{p_1V_1}{p_2} = \frac{250\text{ torr}\cdot2\text{ L}}{100\text{ torr}} = 5" L.


Answer: the helium sample occupies 5 L at 100 torr.